# Direct Imaging of Hydrogen-Driven Dislocation and Strain Field Evolution in a Stainless Steel Grain

### v4

@@ 1 Introduction
Hydrogen is an attractive energy carrier that can be produced through methods such as water electrolysis or photoelectrochemical technologies using renewable resources [1]. This green hydrogen can be used to decarbonize industries that rely heavily on fossil fuels such as aviation [2], heavy transportation [3], industrial production [4], and power generation [1]. As countries strive to reach their Net Zero 2050 sustainability targets, the global production of green hydrogen is forecast to rise annually, reaching 49 million tons per year by 2030 [5]. This necessitates the design of hydrogen fuel systems to store, transport and distribute hydrogen. A long-standing challenge is hydrogen embrittlement (HE), which occurs when hydrogen atoms diffuse into metals, reducing their ductility and strength, leading to premature failure [6]. Austenitic stainless steel (SS), widely used due to its good corrosion resistance and strength, is not immune to this phenomenon. Understanding the mechanisms behind HE is essential to designing the next generation of HE-resistant materials. The study of hydrogen–SS interactions at the nanoscale can provide a mechanistic grounding for multiscale models that seek to predict behavior at larger length scales [7, 8].
@@
HE has been extensively studied at the macroscale through mechanical testing, revealing how cracks initiate and propagate in hydrogen-exposed material [9, 10, 11, 12] and can lead to in-service failure. To explain macroscopic behavior, advances in characterization techniques have enabled the exploration of HE at sub-micron length scales, which has been dominated by electron microscopy (EM) experiments and simulations [13, 14, 15]. This combination of experimental results and theory has led to multiple proposed degradation mechanisms responsible for HE. Here, we will briefly discuss the two most prominent mechanisms. In hydrogen-enhanced localized plasticity (HELP) [9, 10, 16, 17], hydrogen atoms in the metal lattice accumulate at and near the dislocation cores [18] and reduce the elastic stress field, known as hydrogen elastic shielding [19]. This makes it easier for dislocations to move, leading to a localization of deformation, which drives premature material failure. The second mechanism is hydrogen-enhanced decohesion (HEDE) [20, 11], which states that hydrogen atoms accumulate at locations of high triaxial stress, such as grain boundaries or crack tips. The presence of hydrogen reduces the cohesive forces between atoms, weakening the atomic bonds and causing intergranular fracture. The interplay between HELP and HEDE is multifaceted and possibly synergistic [21, 16, 22], though there are still many different proposed theories about the mechanisms involved in HE [16, 23, 24].
@@
Explaining these complex mechanisms requires insights at the nanoscale. Hydrogen can be trapped by vacancies, interstitial sites, solutes, and precipitates, as revealed by thermal desorption spectroscopy (TDS) studies [25]. Hydrogen has been reported to decrease the formation energy of vacancies, leading to an increased vacancy concentration, known as the superabundant vacancies (SAV) model [26]. Vacancies can trap multiple interstitial hydrogen atoms, revealed by positron annihilation spectroscopy (PAS) [27] and density functional theory (DFT) calculations [28]. Though less energetically favored, hydrogen also occupies octahedral interstitial sites in face-centered cubic (FCC) metals [29]. Similarly, hydrogen binds to interstitial and substitutional solutes [27], but this trapping is small compared to vacancy-solute defects [30]. Precipitates such as carbides are also strong hydrogen traps, revealed locally using cryogenic atom probe tomography [18], and some have been shown to increase hydrogen embrittlement resistance by absorbing hydrogen [31].
@@
The ability to detect hydrogen has greatly enhanced our understanding of these mechanisms. Hydrogen microprint techniques [32] and secondary ion mass spectroscopy [33] can provide a static visualization of hydrogen distribution. Scanning Kelvin probe force microscopy [34], neutron tomography [35] and electron-stimulated desorption [36] have been valuable for studying time-resolved hydrogen movement in various microstructures. Despite the availability of these techniques, hydrogen is still generally considered difficult to probe due to its small size, high diffusion rate and low level of interaction with electron probes.
@@
Consequently, there has been a push toward probing in situ responses to hydrogen diffusion to understand degradation mechanisms. Recently, dislocation movement caused by hydrogen desorption has been reported by Koyama et al. [37], where they visualised micron-scale dislocation movements using electron channel contrast imaging (ECCI) on a Fe-Mn–based alloy. Huang et al. [38] revealed bow-out motions of screw dislocations in $\alpha$-Fe pillars undergoing compression tests in a gaseous hydrogen environment inside a transmission electron microscope (TEM). This is an extension to the numerous TEM experiments done by the Robertson group to study hydrogen-induced dislocation movement in foil samples [39, 40, 41, 10, 42]. However, EM techniques only probe relatively thin subsurface volumes of a material ($\lessapprox 100$ nm) [43], and the question of how hydrogen interacts with a single dislocation within the bulk in situ is still experimentally challenging, not well understood [27, 13], and primarily addressed using simulations [15]. Furthermore, hydrogen elastic shielding has only been inferred based on the observations of dislocation pileup [44], but the hydrogen-modified strain fields have not been directly measured experimentally. This information would greatly enhance our understanding of dislocation-hydrogen interactions to accelerate the development of HE-resistant alloys.
@@
Here we address these shortcomings by using in situ Bragg coherent X-ray diffraction imaging (BCDI) [45] to probe the evolution of a single dislocation and its associated strain field during hydrogen charging (Figure 1a), using a bespoke electrochemical flow cell (Supporting Information and Figure LABEL:supp-fig:flowcell, Supporting Information). BCDI is a nondestructive technique that captures three-dimensional (3D) nanoscale images of crystal defects and strain fields [46, 47, 48, 49, 50, 51] by inverting coherent X-ray diffraction patterns [52] from a finite crystalline sample. To produce a sample with grains suitable for BCDI, we use high pressure torsion (HPT) [53] to refine the grain structure of a 316 SS disk (composition shown in Table LABEL:supp-table:composition, Supporting Information). Powder X-ray diffraction data shows that this leads to a deformation-induced martensitic phase transition (Figure LABEL:supp-fig:powder_XRD, Supporting Information). The austenitic phase reappears after ex situ annealing, along with some controlled grain growth and defect relaxation (Experimental Section). The $111$ Bragg peak from an austentitic, FCC grain is measured before and during hydrogen charging at standard ambient temperature and pressure (Experimental Section). The center of the measured grain was close to the center of the disk and located $1.9\pm 0.2\ \mathrm{\mu m}$ beneath the surface (Figure LABEL:supp-fig:Grain_position, Supporting Information). Phase retrieval algorithms are used to recover the real space electron density and phase from the diffraction data (Experimental Section), revealing information about the grain morphology, dislocations, and strain fields, $\varepsilon_{111}$, along the $[111]$ direction associated with the scattering vector, $\mathbf{Q}_{111}$.
@@ 2.1 Bragg peak evolution
After initiating hydrogen charging at 0 h, we observe an increase in the average lattice parameter, $a_{0}$, (Figure 1b) determined using the center position of the Bragg peak (Experimental Section). There are several mechanisms that could explain this increase in $a_{0}$. It is favorable for hydrogen to be trapped in vacancies [27]. However, the relaxation volume for a vacancy in FCC iron is negative and even with six hydrogen atoms per vacancy only becomes slightly positive [28]. Thus, the accumulation of hydrogen-filled vacancies would not explain the substantial lattice swelling we observe. Alternatively, a Frenkel mechanism could be envisaged, where hydrogen prevents Frenkel defects from recombining, and the interstitial of the Frenkel pair (which has large positive relaxation volume) causes the observed swelling. However, this is unlikely because the formation energy of Frenkel defects is large. An energetically favorable explanation would be hydrogen occupying octahedral interstitial sites, which has been reported to cause lattice expansion [54, 29]. Thus, we conclude that the homogeneous lattice strain observed, as a result of hydrogen charging, is dominated by a volumetric strain due to hydrogen occupying interstitial sites. Here, the homogeneous lattice strain, $e_{111}$, is relative to $a_{0}$ at –3.4 h before charging. Note that $e_{111}$ differs from $\varepsilon_{111}$: the latter is relative to the average lattice parameter at time $t$, $a_{0,\ t}$ (Experimental Section) [50]. The volumetric strain is combined with the relaxation volume of hydrogen in austenitic SS to estimate the local hydrogen concentration (Experimental Section), which increases during charging (Figure 1b).
@@ 2.2 Strain at the grain surface
BCDI reconstructions are very sensitive to strain $\left(\approx 10^{-4}\right)$ and highly strained regions scatter outside the detector solid angle, so they appear as missing amplitude in the reconstructions [46]. Relatively disordered grain boundaries with crystal defects in their vicinity appear rough in the reconstructions [55]. This can be seen in Figure 2a, where the morphology is stable during the hydrogen charging, however, a softening of the grain edges is noticeable after 1.9 h (Figure LABEL:supp-fig:all_isosurfaces, Supporting Information). Grain boundaries are strong hydrogen traps ($\approx 17\ \mathrm{kJ\ mol^{-1}}$)[25], so they are expected accumulate hydrogen locally [56] as recently shown by Chen et al. using cryogenic atom probe tomography [18]. As the grain boundaries becomes better organized, here facilitated by hydrogen, lattice distortion near the boundary reduces and hence the apparent surface reconstructed from BCDI evolves. We do not observe any grain boundary mobility due to the presence of new phases, as noticed in palladium thin film grains undergoing a phase transition under hydrogen partial pressures in a previous BCDI study [49].
@@ 2.3 Dislocation dynamics
A phase vortex in a BCDI reconstruction indicates the spatial position of a dislocation core [46], enabling the mapping of 3D dislocations (Experimental Section). Figure 3 shows the dislocation evolution during hydrogen charging. Initially, the grain shows two small dislocations close to the surface, and one large bow-shaped dislocation that extends across the grain, hereafter referred to as the “large dislocation.” The large dislocation remains unchanged before hydrogen charging and up to 1.9 h after the start of charging. To determine its Burgers vector, we assume that it is a shear loop and lies in the $\left(\bar{1}11\right)$ plane, as expected for FCC metals (we confirm the angle between the dislocation glide plane normal and $\mathbf{Q}_{111}$ is 68.9°, less than 2° from the angle between theoretical $\langle 111\rangle$ directions and within error). Thus, we determine the orientation matrix, $\mathbf{UB}$ (Supporting Information). Glide dislocations in the $\left(\bar{1}11\right)$ plane can have three different Burgers vectors $\mathbf{b}=\frac{a_{0}}{2}[110]$, $\frac{a_{0}}{2}[101]$ or $\frac{a_{0}}{2}[0\bar{1}1]$. However, $\mathbf{Q}_{111}\cdot\mathbf{b}\neq 0$ only for $\frac{a_{0}}{2}[110]$ and $\frac{a_{0}}{2}[101]$, meaning that only dislocations with these Burgers vectors will be visible in a $\mathbf{Q}_{111}$ crystal reflection [51]. By comparing the measured strain field of the large dislocation to a 3D elastic model and discrete dislocation dynamics (DDD) simulations, we determine that the large dislocation has Burgers vector $\mathbf{b}=\frac{a_{0}}{2}[110]$ (Supporting Information, Figures LABEL:supp-fig:dislocation_model_slices and LABEL:supp-fig:dislocation_dynamics_models, and Table LABEL:supp-table:error, Supporting Information).
@@
After 1.9 h of hydrogen charging, the large dislocation begins to glide (Figure 3). This is remarkable, as Figure 2 shows little change in surface strain, and hence externally applied stress that could drive plasticity. It suggests that hydrogen either reduces the Peierls barrier for dislocation glide, or unpins the dislocation from obstacles such as alloying elements [57]. Either way, the introduction of hydrogen enables the dislocation to relax, reducing its line length and thus strain energy (Supporting Information and Figure LABEL:supp-fig:dislocation_elastic_energy, Supporting Information) [56]. A DDD simulation (Experimental Section) is setup with the large dislocation in the position observed before hydrogen charging. The ends of the dislocation are pinned at the grain boundary, consistent with our experimental observation, and the system was then allowed to relax under self-stress. The subsequent dislocation evolution, shown in Figure 3, agrees well with that observed experimentally. This confirms that the dislocation glide observed in experiments is relaxation-driven, suggesting hydrogen-enhanced dislocation mobility and/or depinning.
Following the glide motion, after 3.2 h of hydrogen charging, one end of the large dislocation becomes unpinned and climbs out of the glide plane. Subsequently, from 3.9 h onward, the dislocation remains largely unchanged upon further charging. The climb force on a mixed dislocation is the sum of an elastic interaction force and an osmotic force [57]. Figure 2c shows that the average strain on the grain surface changes little, indicating that the contribution of external stresses to dislocation climb will be small. Furthermore, the unpinned dislocation segment is $\approx 200$ nm from the nearest observable small dislocation, meaning that any dislocation–dislocation forces will be minor [19]. Thus, the climb force on the large dislocation will be dominated by the osmotic force, $F_{\mathrm{os}}$. $F_{\mathrm{os}}$, per unit length of a dislocation line is [57],
@@
where $\mathbf{\xi}$ is the unit dislocation line segment direction. Here we observe climb mobility to be much higher than expected for FCC materials [58]. Climb requires diffusion of self-interstitial atoms or vacancies to the dislocation line. Our results suggest hydrogen-facilitated climb, driven by the formation and migration of vacancies as suggested by the SAV model [26, 59]. Furthermore, hydrogen has been demonstrated to enhance metal atom diffusion [26] and lower the vacancy migration free energy barrier in FCC metals [60], which would also drive hydrogen-facilitated climb.
@@
After climb, the climbing segment lies on the $\left(11\bar{1}\right)$ plane. However, since $\mathbf{b}=\frac{a_{0}}{2}[110]$ is not one of the Burgers vectors in this plane, this segment is sessile, explaining the lack of further evolution beyond 3.9 h. This sudden dislocation pinning may result from a sufficiently high hydrogen concentration accumulating at the dislocation core [61], as recently observed in ferritic steel using ECCI [62]. Alternatively, high concentrations of hydrogenated vacancies could lead to dislocation immobilization, as shown in other in situ TEM experiments on initially mobile dislocations in aluminum [63].

### v5

@@ 1 Introduction
Hydrogen is an attractive energy carrier that can be produced through methods such as water electrolysis or photoelectrochemical technologies using renewable resources [Dawood2020]. This green hydrogen can be used to decarbonize industries that rely heavily on fossil fuels such as aviation [Tiwari2024], heavy transportation [Oliveira2021], industrial production [Griffiths2021], and power generation [Dawood2020]. As countries strive to reach their Net Zero 2050 sustainability targets, the global production of green hydrogen is forecast to rise annually, reaching 49 million tons per year by 2030 [IEA2024]. This necessitates the design of hydrogen fuel systems to store, transport and distribute hydrogen. A long-standing challenge is hydrogen embrittlement (HE), which occurs when hydrogen atoms diffuse into metals, reducing their ductility and strength, leading to premature failure [Johnson1875]. Austenitic stainless steel (SS), widely used due to its good corrosion resistance and strength, is not immune to this phenomenon. Understanding the mechanisms behind HE is essential to designing the next generation of HE-resistant materials. The study of hydrogen–SS interactions at the nanoscale can provide a mechanistic grounding for multiscale models that seek to predict behavior at larger length scales [Castelluccio2018, Li2015b].
@@
HE has been extensively studied at the macroscale through mechanical testing, revealing how cracks initiate and propagate in hydrogen-exposed material [Beachem1972, Shih1988, Oriani1974, Lynch1988] and can lead to in-service failure. To explain macroscopic behavior, advances in characterization techniques have enabled the exploration of HE at sub-micron length scales, which has been dominated by electron microscopy (EM) experiments and simulations [Martin2019, Chen2024a, Dong2022]. This combination of experimental results and theory has led to multiple proposed degradation mechanisms responsible for HE. Here, we will briefly discuss the two most prominent mechanisms. In hydrogen-enhanced localized plasticity (HELP) [Beachem1972, Shih1988, Robertson2015, Robertson2009], hydrogen atoms in the metal lattice accumulate at and near the dislocation cores [Chen2020a] and reduce the elastic stress field, known as hydrogen elastic shielding [Birnbaum1994]. This makes it easier for dislocations to move, leading to a localization of deformation, which drives premature material failure. The second mechanism is hydrogen-enhanced decohesion (HEDE) [Pfeil1926, Oriani1974], which states that hydrogen atoms accumulate at locations of high triaxial stress, such as grain boundaries or crack tips. The presence of hydrogen reduces the cohesive forces between atoms, weakening the atomic bonds and causing intergranular fracture. The interplay between HELP and HEDE is multifaceted and possibly synergistic [Djukic2019, Robertson2015, Koyama2014], though there are still many different proposed theories about the mechanisms involved in HE [Robertson2015, Lynch2012, Nagumo2004].
@@
Explaining these complex mechanisms requires insights at the nanoscale. Hydrogen can be trapped by vacancies, interstitial sites, solutes, and precipitates, as revealed by thermal desorption spectroscopy (TDS) studies [Choo1982a]. Hydrogen has been reported to decrease the formation energy of vacancies, leading to an increased vacancy concentration, known as the superabundant vacancies (SAV) model [Fukai2003]. Vacancies can trap multiple interstitial hydrogen atoms, revealed by positron annihilation spectroscopy (PAS) [Myers1992] and density functional theory (DFT) calculations [Nazarov2010]. Though less energetically favored, hydrogen also occupies octahedral interstitial sites in face-centered cubic (FCC) metals [Fukai2005]. Similarly, hydrogen binds to interstitial and substitutional solutes [Myers1992], but this trapping is small compared to vacancy-solute defects [Li2023]. Precipitates such as carbides are also strong hydrogen traps, revealed locally using cryogenic atom probe tomography [Chen2020a], and some have been shown to increase hydrogen embrittlement resistance by absorbing hydrogen [Bhadeshia2016].
@@
The ability to detect hydrogen has greatly enhanced our understanding of these mechanisms. Hydrogen microprint techniques [Ovejero-García1985] and secondary ion mass spectroscopy [Takai1995] can provide a static visualization of hydrogen distribution. Scanning Kelvin probe force microscopy [Evers2013], neutron tomography [Griesche2014] and electron-stimulated desorption [Miyauchi2020] have been valuable for studying time-resolved hydrogen movement in various microstructures. Despite the availability of these techniques, hydrogen is still generally considered difficult to probe due to its small size, high diffusion rate and low level of interaction with electron probes.
@@
Consequently, there has been a push toward probing in situ responses to hydrogen diffusion to understand degradation mechanisms. Recently, dislocation movement caused by hydrogen desorption has been reported by Koyama et al. [Koyama2020], where they visualised micron-scale dislocation movements using electron channel contrast imaging (ECCI) on a Fe-Mn–based alloy. Huang et al. [Huang2023] revealed bow-out motions of screw dislocations in $\alpha$-Fe pillars undergoing compression tests in a gaseous hydrogen environment inside a transmission electron microscope (TEM). This is an extension to the numerous TEM experiments done by the Robertson group to study hydrogen-induced dislocation movement in foil samples [Robertson2001, Sofronis2002, Robertson1984, Shih1988, Wang2014]. However, EM techniques only probe relatively thin subsurface volumes of a material ($\lessapprox 100$ nm) [Williams2009], and the question of how hydrogen interacts with a single dislocation within the bulk in situ is still experimentally challenging, not well understood [Myers1992, Martin2019], and primarily addressed using simulations [Dong2022]. Furthermore, hydrogen elastic shielding has only been inferred based on the observations of dislocation pileup [Ferreira1998], but the hydrogen-modified strain fields have not been directly measured experimentally. This information would greatly enhance our understanding of dislocation-hydrogen interactions to accelerate the development of HE-resistant alloys.
@@
Here we address these shortcomings by using in situ Bragg coherent X-ray diffraction imaging (BCDI) [Pfeifer2006] to probe the evolution of a single dislocation and its associated strain field during hydrogen charging (Figure 1a), using a bespoke electrochemical flow cell (Supporting Information and Figure LABEL:supp-fig:flowcell, Supporting Information). BCDI is a nondestructive technique that captures three-dimensional (3D) nanoscale images of crystal defects and strain fields [Clark2015, Orr2023, Singer2018, Yau2017b, Atlan2023a, Hofmann2020] by inverting coherent X-ray diffraction patterns [Fienup1982] from a finite crystalline sample. To produce a sample with grains suitable for BCDI, we use high pressure torsion (HPT) [Zhilyaev2008] to refine the grain structure of a 316 SS disk (composition shown in Table LABEL:supp-table:composition, Supporting Information). Powder X-ray diffraction data shows that this leads to a deformation-induced martensitic phase transition (Figure LABEL:supp-fig:powder_XRD, Supporting Information). The austenitic phase reappears after ex situ annealing, along with some controlled grain growth and defect relaxation (Experimental Section). The $111$ Bragg peak from an austentitic, FCC grain is measured before and during hydrogen charging at standard ambient temperature and pressure (Experimental Section). The center of the measured grain was close to the center of the disk and located $1.9\pm 0.2\ \mathrm{\mu m}$ beneath the surface (Figure LABEL:supp-fig:Grain_position, Supporting Information). Phase retrieval algorithms are used to recover the real space electron density and phase from the diffraction data (Experimental Section), revealing information about the grain morphology, dislocations, and strain fields, $\varepsilon_{111}$, along the $[111]$ direction associated with the scattering vector, $\mathbf{Q}_{111}$.
@@ 2.1 Bragg peak evolution
After initiating hydrogen charging at 0 h, we observe an increase in the average lattice parameter, $a_{0}$, (Figure 1b) determined using the center position of the Bragg peak (Experimental Section). There are several mechanisms that could explain this increase in $a_{0}$. It is favorable for hydrogen to be trapped in vacancies [Myers1992]. However, the relaxation volume for a vacancy in FCC iron is negative and even with six hydrogen atoms per vacancy only becomes slightly positive [Nazarov2010]. Thus, the accumulation of hydrogen-filled vacancies would not explain the substantial lattice swelling we observe. Alternatively, a Frenkel mechanism could be envisaged, where hydrogen prevents Frenkel defects from recombining, and the interstitial of the Frenkel pair (which has large positive relaxation volume) causes the observed swelling. However, this is unlikely because the formation energy of Frenkel defects is large. An energetically favorable explanation would be hydrogen occupying octahedral interstitial sites, which has been reported to cause lattice expansion [Ulmer1993, Fukai2005]. Thus, we conclude that the homogeneous lattice strain observed, as a result of hydrogen charging, is dominated by a volumetric strain due to hydrogen occupying interstitial sites. Here, the homogeneous lattice strain, $e_{111}$, is relative to $a_{0}$ at –3.4 h before charging. Note that $e_{111}$ differs from $\varepsilon_{111}$: the latter is relative to the average lattice parameter at time $t$, $a_{0,\ t}$ (Experimental Section) [Atlan2023a]. The volumetric strain is combined with the relaxation volume of hydrogen in austenitic SS to estimate the local hydrogen concentration (Experimental Section), which increases during charging (Figure 1b).
@@ 2.2 Strain at the grain surface
BCDI reconstructions are very sensitive to strain $\left(\approx 10^{-4}\right)$ and highly strained regions scatter outside the detector solid angle, so they appear as missing amplitude in the reconstructions [Clark2015]. Relatively disordered grain boundaries with crystal defects in their vicinity appear rough in the reconstructions [Yang2021]. This can be seen in Figure 2a, where the morphology is stable during the hydrogen charging, however, a softening of the grain edges is noticeable after 1.9 h (Figure LABEL:supp-fig:all_isosurfaces, Supporting Information). Grain boundaries are strong hydrogen traps ($\approx 17\ \mathrm{kJ\ mol^{-1}}$)[Choo1982a], so they are expected accumulate hydrogen locally [Kirchheim2007a] as recently shown by Chen et al. using cryogenic atom probe tomography [Chen2020a]. As the grain boundaries becomes better organized, here facilitated by hydrogen, lattice distortion near the boundary reduces and hence the apparent surface reconstructed from BCDI evolves. We do not observe any grain boundary mobility due to the presence of new phases, as noticed in palladium thin film grains undergoing a phase transition under hydrogen partial pressures in a previous BCDI study [Yau2017b].
@@ 2.3 Dislocation dynamics
A phase vortex in a BCDI reconstruction indicates the spatial position of a dislocation core [Clark2015], enabling the mapping of 3D dislocations (Experimental Section). Figure 3 shows the dislocation evolution during hydrogen charging. Initially, the grain shows two small dislocations close to the surface, and one large bow-shaped dislocation that extends across the grain, hereafter referred to as the “large dislocation.” The large dislocation remains unchanged before hydrogen charging and up to 1.9 h after the start of charging. To determine its Burgers vector, we assume that it is a shear loop and lies in the $\left(\bar{1}11\right)$ plane, as expected for FCC metals (we confirm the angle between the dislocation glide plane normal and $\mathbf{Q}_{111}$ is 68.9°, less than 2° from the angle between theoretical $\langle 111\rangle$ directions and within error). Thus, we determine the orientation matrix, $\mathbf{UB}$ (Supporting Information). Glide dislocations in the $\left(\bar{1}11\right)$ plane can have three different Burgers vectors $\mathbf{b}=\frac{a_{0}}{2}[110]$, $\frac{a_{0}}{2}[101]$ or $\frac{a_{0}}{2}[0\bar{1}1]$. However, $\mathbf{Q}_{111}\cdot\mathbf{b}\neq 0$ only for $\frac{a_{0}}{2}[110]$ and $\frac{a_{0}}{2}[101]$, meaning that only dislocations with these Burgers vectors will be visible in a $\mathbf{Q}_{111}$ crystal reflection [Hofmann2020]. By comparing the measured strain field of the large dislocation to a 3D elastic model and discrete dislocation dynamics (DDD) simulations, we determine that the large dislocation has Burgers vector $\mathbf{b}=\frac{a_{0}}{2}[110]$ (Supporting Information, Figures LABEL:supp-fig:dislocation_model_slices and LABEL:supp-fig:dislocation_dynamics_models, and Table LABEL:supp-table:error, Supporting Information).
@@
After 1.9 h of hydrogen charging, the large dislocation begins to glide (Figure 3). This is remarkable, as Figure 2 shows little change in surface strain, and hence externally applied stress that could drive plasticity. It suggests that hydrogen either reduces the Peierls barrier for dislocation glide, or unpins the dislocation from obstacles such as alloying elements [Anderson2017]. Either way, the introduction of hydrogen enables the dislocation to relax, reducing its line length and thus strain energy (Supporting Information and Figure LABEL:supp-fig:dislocation_elastic_energy, Supporting Information) [Kirchheim2007a]. A DDD simulation (Experimental Section) is setup with the large dislocation in the position observed before hydrogen charging. The ends of the dislocation are pinned at the grain boundary, consistent with our experimental observation, and the system was then allowed to relax under self-stress. The subsequent dislocation evolution, shown in Figure 3, agrees well with that observed experimentally. This confirms that the dislocation glide observed in experiments is relaxation-driven, suggesting hydrogen-enhanced dislocation mobility and/or depinning.
Following the glide motion, after 3.2 h of hydrogen charging, one end of the large dislocation becomes unpinned and climbs out of the glide plane. Subsequently, from 3.9 h onward, the dislocation remains largely unchanged upon further charging. The climb force on a mixed dislocation is the sum of an elastic interaction force and an osmotic force [Anderson2017]. Figure 2c shows that the average strain on the grain surface changes little, indicating that the contribution of external stresses to dislocation climb will be small. Furthermore, the unpinned dislocation segment is $\approx 200$ nm from the nearest observable small dislocation, meaning that any dislocation–dislocation forces will be minor [Birnbaum1994]. Thus, the climb force on the large dislocation will be dominated by the osmotic force, $F_{\mathrm{os}}$. $F_{\mathrm{os}}$, per unit length of a dislocation line is [Anderson2017],
@@
where $\mathbf{\xi}$ is the unit dislocation line segment direction. Here we observe climb mobility to be much higher than expected for FCC materials [Abu-Odeh2020]. Climb requires diffusion of self-interstitial atoms or vacancies to the dislocation line. Our results suggest hydrogen-facilitated climb, driven by the formation and migration of vacancies as suggested by the SAV model [Fukai2003, Nagumo2001]. Furthermore, hydrogen has been demonstrated to enhance metal atom diffusion [Fukai2003] and lower the vacancy migration free energy barrier in FCC metals [Du2020], which would also drive hydrogen-facilitated climb.
@@
After climb, the climbing segment lies on the $\left(11\bar{1}\right)$ plane. However, since $\mathbf{b}=\frac{a_{0}}{2}[110]$ is not one of the Burgers vectors in this plane, this segment is sessile, explaining the lack of further evolution beyond 3.9 h. This sudden dislocation pinning may result from a sufficiently high hydrogen concentration accumulating at the dislocation core [Murakami2010], as recently observed in ferritic steel using ECCI [Kim2025]. Alternatively, high concentrations of hydrogenated vacancies could lead to dislocation immobilization, as shown in other in situ TEM experiments on initially mobile dislocations in aluminum [Xie2016].